OpenStudy (anonymous):

Write the expression (ln) 3x^2/(x+3)^5 in an equivalent form that has no powers, products, or quotients in it.

OpenStudy (anonymous):

=\[\ln 3x ^{2}-\ln (x+3)^{5}=2\ln 3x-5\ln (x+3)\]

OpenStudy (anonymous):

is that good for you?